Subject: 3.5-m Users Committee minutes 7/16/07
From: Bruce Gillespie
Submitted: Fri, 20 Jul 2007 14:23:56 -0600
Message number: 1043
(previous: 1042,
next: 1044
up: Index)
APO 3.5-m Users Committee Phonecon, 7/16/07
Attending: Suzanne Hawley, Bruce Gillespie, Mark Klaene, Scott
Anderson, Remy Indebetouw, Al Harper, Michael Strauss, Russet
McMillan, Jon Fulbright, John Bally, Grace Wolf-Chase, Bill Ketzeback
Absent: Jon Holtzman
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael had nothing to report, except his hopes
for improved observing weather. Mark said that the short-term
weather forecast is predicting better weather.
CU (Bally) - John had nothing to report from the users. There are
issues regarding the upcoming NIC-FPS upgrades, namely a potential
staffing shortfall at Boulder. A meeting at CU will be held this
afternoon to identify people and funding. The IR test dewar for
testing the Fabry-Perot has arrived and is being tested. There is a
risk that certain key staff will not be available to work on the
instrument during the shutdown, and finishing the instrument
documentation is also a concern. Mark Klaene mentioned that with the
16-channel read upgrade to the detector, we may experience file-size
and site-to-user bandwidth problems. We may need to look into data
compression and/or next-day raw data retrieval. A reminder to users:
NIC-FPS will be taken off line beginning 26 July for the upgrade work.
NMSU (Holtzman) - No report available.
Washington (Anderson) - Scott had nothing to report. Suzanne added
that she is currently teaching a grad course on observing, and has
found that the site user instrumentation documentation is still an
issue. The DIS documentation is great, but the other instruments'
documentation needs improvement. She also mentioned that if other
institutions have observing courses using APO, they should let her
know how those are going and if there are any issues that the
observatory should address
JHU (Fulbright) - Jon said that the recent replacement of the echelle
calibration lamp has worked well. Suzanne asked if JHU and Chicago
hold observing classes. Al said that UC has an experimental
astronomy course that largely deals with instrumentation, not
observing, and he would suggest adding an observing course to the
curriculum committee. Jon said that JHU treats observer training on
an individual basis. Michael said that at Princeton most of the
observers are post-docs who get APO-specific training individually at
the site.
Virginia (Indebetouw) - Remy reported that the UVa users are pleased
with their data. They have had installation problems with TUI
related to firewalls and libraries [note: institutional system
administrators should contact Russell Owen at UW--
rowen@u.washington.edu--for advice on TUI installations]. Remy also
said that his users have tried using the new SPIcam/TUI software with
mixed results. Russet said that there is still testing to be done
that has been thwarted by recent cloudy weather.
Chicago (Harper) - Al had nothing to add that wasn't already
discussed above.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below; Mark walked through the highlights. The
so-called "monsoon" storms have arrived. Jim Davenport is now
soloing as a summer fill-in Observing Specialist before he heads off
to grad school next month. Bill Ketzeback is back at work, and Jack
Dembicky is on family-related leave.
The only new thing to report on the telescope is a recent glitch with
tertiary mirror rotation, which was discovered last week. The
rotation from port to port seems to have been become jerky and
intermittent. A loose wire on the rotation drive brake system was
discovered and repaired this morning, and subsequent tests so far
have shown no motion failures. Further testing will be done, and
Russell and Mark will devise a software
change to limit the acceleration/speed of the tertiary rotation to
protect the mechanicals if this anomaly repeats. Mark reminded us
that the summer shutdown is still scheduled from 20 August through 16
September. There is only limited on-sky testing time available at
the end of the shutdown, and we are counting on favorable weather in
the last few days of the shutdown to re-commission the telescope.
Mark reported that Agile is about to have another commissioning run
at the telescope. Also, the DIS high-red replacement grating is
being tested at JHU prior to delivery to APO and installation. The
Lakeshore controller that controls the temperature of the DIS red
chip has failed and is under repair. In the meantime users should
make sure to take concurrent flats and biases with their DIS-red
data, as these may be slightly different than before the Lakeshore
failed.
For the CIF projects, the instrument rotator parts are in, and Fritz
Stauffer and Nick McDonald are visiting engineers at KPNO and
Flagstaff to discuss various axis controller and motor questions.
For the rebuild of the tertiary mirror mount, we still need to hear
from users on requirements for tip-tilt so we can plan the scope of
the tertiary mount redesign. Suzanne said we need to hear from users
what, if any, are the compelling scientific requirements for tip-tilt
are, and what the derived engineering requirements are. Jon Bally
said that the CU users want the hooks left in for tip-tilt in the
tertiary. Remy asked who do we talk to about tip-tilt.
ACTION: Suzanne said that the Users Committee members need to
identify which of their users would be willing to contribute effort
and participate in a telecon to scope tip-tilt for the 3.5-m, and to
communicate their names to her and Bruce Gillespie before 30 July.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 6/10/07
through 7/5/07
Mark Klaene
0) Overview
Monsoonal weather started a little early this year with moderate rain
and evening clouds limiting observations for much of the past 2
weeks. The June engineering time was clouded out; however, some in-
dome testing of the SPIcam TUI interface and further training of Jim
Davenport made the time useful. Bill Ketzeback has returned from his
family medical leave.
1) Telescope
No significant problems this period. Tracking tuning was not
accomplished, but tracking has generally been acceptable, not
requiring that we pre-empt observing.
The summer shutdown is still on track for August 20 to September 16.
The driver is the primary mirror re-aluminization. We have a limited
number of days on the sky for recovery; hopefully the weather will
cooperate and allow us to return to science on schedule.
2) Instruments
Agile is preparing for more commissioning runs at the telescope this
month. The fiber interface was working when it left Princeton
Instruments. The focal reducer is planned to be completed and ready
for commissioning during July.
The replacement DIS Red high-resolution grating has been received at
JHU for testing. Delivery to the observatory and installation in DIS
is expected later this month.
The Lakeshore temperature controller failed on both DIS cameras--
repair is in work. This leaves us with no DIS red heater and only
the ion pump controllers as a way to gauge temp and vacuum.
We replaced the ThAr lamp on the Echelle.
Incorporation of SPIcam into TUI has been released to users. Some
scripts are still under development.
NA2 guider camera had an intermittent network connection failure
develop. The cause was traced to a broken network cable. A more
flexible cable has been installed.
NIC-FPS is scheduled to be taken off line and sent to CU on the 26th
of July. Work will include implementing 16-channel readout capability
and F-P Etalon installation and testing.
3) Engineering and CIF projects
2007:
Rotator upgrades: two machining work orders have been let and the PO
for the motor and accessories has been signed. Expect delivery late
August.
Axis controllers: bench setup has been built and is undergoing
testing of the new axis controllers.
2008:
Tertiary mounting: discussions are ongoing as to the need for tip/
tilt to be incorporated. Input from the scientific community is
required to determine if fast guiding at the tertiary is desired. If
interested please let Larry Carey know (lcarey@astro.washington.edu)
Direct drives: research into direct drives for the azimuth and
altitude drives continues.
4) Miscellaneous
Jim Davenport completed his training as a summer fill-in Observing
Specialist. Jim will be soloing during the rest of July and August.
**********************************
SPIcam/TUI interface - Klaene:
The on-sky testing of the new SPIcam/TUI interface has been hampered
by bad weather during engineering time. The readout speed is an
issue, which is being worked on by Craig Loomis at PU. Russet said
that users have been trying it, and bugs are being chased down and
tested, but SPIcam is not a frequently scheduled instrument. There
are operational workarounds that speed things up, but they need
further real-life tests. Suzanne said we greatly appreciate users
who are willing to help us test and debug SPIcam/TUI on a shared-risk
basis, but users can still use the old system if they prefer. We
will continue to work on making the new software the default, but it
will depend on how much on-sky testing time we can get over the
coming weeks.
**********************************
Management changes at APO - Hawley:
Suzanne pointed to the recent announcement that Bruce Gillespie has
been reassigned to the position of ARC Program Administrator, in
which he will act as the AS2 Program Manager and also provide part-
time support to her and Rich Kron for the 3.5-m telescope and SDSS-
II, respectively. He will continue his role with 3.5-m Users
Committee and its constituent institutions through a transitional
period during the coming year. Mark Klaene has been promoted to the
position of APO Site Operations Manager, and will be increasingly
involved with 3.5-m planning and the user community.
**********************************
Actions from earlier meetings - group
[open from previous months]:
none
[new actions from this month]:
ACTION: Suzanne said that the Users Committee members need to
identify which of their users would be willing to contribute effort
and participate in a telecon to scope tip-tilt for the 3.5-m, and to
communicate their names to her and Bruce Gillespie before 30 July.
**********************************
[Suzanne asked me to post this addendum to the minutes]:
To members of the 3.5m users community:
As documented in recent 3.5m users committee minutes, the NSF has
charged a committee (called ReSTAR) with carrying out a survey of the
astronomical community to assess the interest in supporting
scientific programs on small and moderate sized (less than 6m)
telescopes. This comes from the Senior Review report, and the 3rd
Telescope Systems Workshop report, both of which highlighted the need
for telescopes of this size. Obviously the 3.5m falls into this
class. If you have not already done so, please take a moment to go
to the ReSTAR web page and provide your input. An NSF program that
provides funding for telescopes of this size could be an attractive
option for improving our instrumentation and supporting our
operations in the future. The committee web page is located at:
http://www.noao.edu/system/restar
and the form to submit input is linked off that page. The committee
will meet next on July 30-31 to consider the input; responses
received before then will be most likely to be effective.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 17 September, at
8:30 AM Pacific Time (There will be no phonecon in August). The
agenda and other materials will be sent to the committee members
during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1043 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO